OpenStudy (anonymous):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder?

OpenStudy (anonymous):

Wouldn't I divide the volume by three and get 8?

OpenStudy (perl):

we know that the height of the cone and the radius of the cone must be the same as the height of the cylinder and the radius of the cylinder (since it fits in exactly)

OpenStudy (perl):



OpenStudy (perl):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? 24 = π r^2 * h V = 1/3 π r^2 * h = 1/3 ( 24) yes you are correct

OpenStudy (anonymous):

thanks :)